1)IF sinx=cos 2 x, then cos 2 x(1 + cos 2 x) is? 2)SOLUTION SET OF (3/5) x =(x -x 2 -9) is?

1)IF sinx=cos2x, then cos2x(1 + cos2x) is?


2)SOLUTION SET OF (3/5)x=(x -x2-9) is?

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear vardaan

sinx=cos2x

sinx =1-sin2x

sin2x = 1-sinx  ..............1

 

now   cos2x(1 + cos2x) = sinx (1+sinx)

                                            =sinx + sin2x

                                            = sinx + 1-sinx         from equation (1)

                                             =1

 

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin